OpenStudy (anonymous):

The volume of a rectangular box is limited to 1100 cu. in. If the height is 8 inches and the width is 10 inches how long can the box be?

OpenStudy (ghazi):

l*b*h=1100 therefore 10*8*x=1100 now solve for x

OpenStudy (anonymous):

V=1100 H=8in B=10in L=? V=l*b*h 1100=l*8*10 1100=80l 1100/80=l 13.75=l Length of box id 13.75in

OpenStudy (ghazi):

yes